# Efficient Decoder for End-to-End Oriented Object Detection in Remote Sensing Images

### v1

@@ IV-E Ablation Study
Figure 6: Some detection results on DOTA-v1.0. The first row of images shows the detection results. The second row of images shows RRoI attention.
@@
IV-F visualization
Visualization of queries through Selective Distinct Queris. Shown in Fig. 5, we visualize initial queries, selective queries, distinct queries, and final queries. Initial queries are more random. Selective queries are more accurate than initial queries. We filter similar queries to obtain distinct queries. Final queries can detect objects accurately.
Visualization of RRoI attention We visualize RRoI attention in Fig. 6 without using CAM [57]. We observe that it mostly focuses on object extremities such as the front and back of planes and ships. These highlighted areas are very similar to DETR [6] and Deformable DETR [7].
Visualization of results We visualize some predicted results in Fig. 6 and Fig. 7. We can see that our model detects objects accurately.
V Conclusions
In this paper, we propose an end-to-end detector for oriented object detection in remote sensing images, including two key technologies RRoI attention and SDQ. The proposed RRoI attention align image features and oriented boxes through a cross-attention mechanism for objects with arbitrary orientations and varying scales. To avoid a large number of initial queries or extra auxiliary branches and eliminate the similar queries hindering the optimization, the SDQ first collects high-quality selective queries from intermediate decoder layers, and second filters similar queries to keep remaining queries distinct which facilitates the one-to-one label assignment. Extensive experiments show that our method has competitive accuracy to the modern CNN-based one-stage and two-stage detectors and transformer-based detectors. We hope it will inspire researchers to further explore the application of end-to-end models in oriented object detection.

### v4

@@ 4.8 Effects of Queries
Figure 9: Some detection results on DOTA-v1.0. The left column of images shows the detection results. The right column of images shows RRoI attention.
4.9 Comparison of similar queries
We calculate the ratios of similar queries for our method and other end-to-end methods on DOTA-v1.0 to further validate the effectiveness of SDQ. The formula is shown as Eq. 9 and results are illustrated in Fig. 6. Our approach effectively filters out similar queries, resulting in a notable reduction in their ratio across four different IoU threshold settings: 0.8, 0.85, 0.9, and 0.95. Taking a threshold of 0.95 as an example, the ratios of similar queries are 15.20%, 9.63%, 10.16%, and 3.86% in Deformable DETR-O, Deformable DETR-O with CSL, ARS-DETR, and DDQ-O, respectively. Our method can significantly reduce it to 0.79%, which serves as evidence of its effectiveness.
4.10 Convergence
DETR (Carion et al., 2020) suffers from long training time and slow convergence. The Deformable DETR series (Zhu et al., 2020; Zeng et al., 2024) methods alleviate this issue. To further investigate training efficiency, we compare RQFormer with other end-to-end models. Convergence curves for these models are depicted in Fig. 7. For a fair comparison, all methods are trained on 12 epochs on DIOR-R using ResNet50 with 300 queries. The learning rate decays by 10$\times$ on the $8^{th}$ and $11^{th}$ epochs respectively. RQFormer achieves a mAP of 62.8% in just 12 epochs, surpassing Deformable DETR-O with CSL (Yang and Yan, 2020) (31.23%) and ARS-DETR (Zeng et al., 2024) (38.9%). Additionally, RQFormer outperforms baseline DDQ-O by 1.1%. These results underscore the effectiveness of our approach in mitigating the challenges associated with slow convergence.
4.11 Visualization
Visualization of queries through Selective Distinct Queries. Shown in Fig. 8, we visualize initial queries, selective queries, distinct queries, and final queries. Initial queries are more random. There are many highly overlapping boxes. Selective queries are more accurate than initial queries. They effectively cover objects. We filter similar queries to obtain distinct queries. Boxes with high overlapping are filtered out. Final queries can detect objects accurately.
(a) RQFormers(ours)
(b) ARS-DETR
(c) OrientedFormer
(d) Oriented R-CNN
(e) SASM
Figure 10: Comparison between our method and others. The confidence threshold is set to 0.3 for better visualization. The 1st line: objects with large aspect ratio, e.g., harbor. The 2nd line: densely packed objects, e.g., planes. The 3rd line: square-like objects, e.g., baseball diamond.
@@ 4.11 Visualization
Figure 12: Example results on scene text datasets ICDAR2015.
Visualization of RRoI attention. We visualize RRoI attention in Fig. 9 without using CAM. We observe that it mostly focuses on object extremities such as the front and back of planes and ships. These highlighted areas are very similar to DETR (Carion et al., 2020) and Deformable DETR (Zhu et al., 2020).
Comparison with other methods. We compare our method with others on large aspect ratio objects, densely packed objects, and square-like objects, as shown in Fig. 10. Other methods failed to detect these challenging objects.
Visualization of results. We visualize some predicted results on DOTA-v1.0 in Fig. 9, DIOR-R in Fig. 11, and ICDAR2015 in Fig. 12. We can see that our model detects objects accurately both in positions and categories.
5 Conclusions
In this paper, we propose an end-to-end detector RQFormer for oriented object detection that incorporates two key technologies: Rotated RoI Attention (RRoI Attention) and Selective Distinct Queries (SDQ). The proposed RRoI Attention mechanism aligns keys and positional queries through cross-attention, effectively handling objects with arbitrary orientations and varying scales. To prevent similar queries from hindering optimization, SDQ first collects high-quality selective queries from intermediate decoder layers and then filters out similar ones, ensuring that the remaining queries are distinct. This process facilitates one-to-one label assignment. Extensive experiments demonstrate that our method achieves competitive accuracy compared to modern CNN-based one-stage and two-stage detectors, as well as transformer-based detectors. We hope our approach will inspire researchers to further explore the application of end-to-end models in oriented object detection.
Limitations and feature works. The Rotated RoI attention primarily focuses on local areas strictly within positional queries. Contextual information around these boxes may enhance attention. Further exploration into integrating contextual information into attention is necessary.
CRediT authorship contribution statement
Jiaqi Zhao: Conceptualization, Methodology, Software, Writing – original draft, Writing – review & editing. Zeyu Ding: Conceptualization, Methodology, Software, Writing – original draft, Writing – review & editing. Yong Zhou: Visualization, Formal analysis, Funding acquisition, Supervision. Hancheng Zhu: Data curation, Investigation, Writing – review & editing. Wen-Liang Du: Data curation, Investigation, Writing – review & editing. Rui Yao: Software, Validation. Abdulmotaleb El Saddik: Validation, Writing – review & editing.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
Data will be made available on request.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (No.62272461,62002360, 62172417, 62101555), the Double FirstClass Project of China University of Mining and Technology for Independent Innovation and Social Service under Grant 2022ZZCX06, the Six Talent Peaks Project in Jiangsu Province (No. 2015-DZXX-010, 2018-XYDXX-044).